Log 2 = 0.301 log 3 = 0.477 log 5 = 0.699 using this information
Find:- a) Log 30 b)Log o.6

Answer by ankor@dixie-net.com(22740)

Log 2 = 0.301 log 3 = 0.477 log 5 = 0.699 using this information
Find:- a) Log 30 b)Log o.6
:
a)
We know: 2 * 3 * 5 = 30
therefore: log(2) + log(3) + log(5) = log(30)
then
adding the logs: .301 + . 477 + .699 = 1.477 is log of 30
:
b)
We know: 2 * 3 * .1 = .6 and we also know the log of .1 = -1
therefore: log(2) + log(3) + log(.1) = log(.6)
adding this: .301 + .477 - 1 = -.222 is the log of .1